OpenStudy (anonymous):

Solve the following system of equations: x + 3y - z = 2 x - 2y + 3z = 7 x + 2y - 5z = -21

OpenStudy (igreen):

@kelseymanghamx Any options?

OpenStudy (anonymous):

Yes. Sorry. @iGreen. (2, 3, 5) (-2, 3, 5) (2, -3, 5) (2, 3, -5)

OpenStudy (igreen):

Okay, I never really learned this before, but we can plug the solutions into the three equations to find out our answer.

OpenStudy (igreen):

\(x + 3y - z = 2\) \((2, 3, 5)\) \(2 + 3(3) - 5 = 2\) \(2 + 9 - 5 = 2\) \(11 - 5 = 2\) \(6 = 2\) 6 does NOT equal 2. So the first option is wrong.

OpenStudy (anonymous):

I am not a strong math person, and I am trying to learn online. I'll help as best as I can. How do you know where to plug in the 3 numbers.

OpenStudy (igreen):

\(x + 3y - z = 2\) \( (-2, 3, 5)\) \(-2 + 3(3) - 5 = 2\) \(-2 + 9 - 5 = 2\) \( 7 - 5 = 2\) \( 2 = 2\) This is true, now let's go to the second equation: \(x - 2y + 3z = 7\) \( (-2, 3, 5)\) \(-2 - 2(3) + 3(5) = 7\) \(-2 - 6 + 3(5) = 7\) \(-2 - 6 + 15 = 7\) \(-8 + 15 = 7\) \(8 = 7\) This is FALSE, so we know Option B is wrong too.

OpenStudy (igreen):

The numbers in the options are in the order of (x, y, z). So look in the equation for x, y, and z then plug them in.

OpenStudy (igreen):

\(x + 3y - z = 2\) \( (2, -3, 5)\) \(2 + 3(-3) - 5 = 2\) \(2 - 9 - 5 = 2\) \(-7 - 5 = 2\) \(-12 = 2\) This is also false, which leaves you with Option D.

OpenStudy (anonymous):

Thank you so much! This is super helpful.

OpenStudy (igreen):

Sorry, I made a mistake on Option B.

OpenStudy (igreen):

Which is your answer..

OpenStudy (anonymous):

Option B is my answer?